OpenStudy (anonymous):

How do you solve 5x² - 20x² (GCF)

OpenStudy (aravindg):

GCF means?

OpenStudy (aravindg):

-15x^2

OpenStudy (anonymous):

GCF means Greatest Common Factor.

OpenStudy (anonymous):

Am I right?

OpenStudy (aravindg):

@allea_marie reply pls

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

There needs to be more here, for there to be something to solve.

OpenStudy (anonymous):

Yes it stands for Greatest Common Factor

OpenStudy (anonymous):

so what are the two numbers that you want to find the greatest common factor of?

OpenStudy (anonymous):

But why do you need GCF to solve this? It is quite simple. Just subtract 5-20 and carry over the x^2. That would be -15x^2

OpenStudy (anonymous):

Okay thanks, how would you solve 2x² - 21x=54

OpenStudy (anonymous):

that one can be solved by factorising the polynomial.

OpenStudy (anonymous):

\[ 2x^2 - 21x =54 \] \[2x^2 - 21x - 54 = 0 \] Now you want to solve: \[(2x + a)(x +b) = 2x^2 - 21x - 54 \] Note that you must have \[ a*b = -54 \]

OpenStudy (anonymous):

this won't always work so easily. But I can see here that it does.

OpenStudy (anonymous):

Okay thank you